OpenStudy (anonymous):

greatest common factor of each monomial 46a^2 28a 20a^6

OpenStudy (anonymous):

of each or of all?

OpenStudy (campbell_st):

look at it this way \[46a^2 = 2a \times 23a\] \[28a = 2a \times 14\] \[20a^2 = 2a \times 10a\] this may help

OpenStudy (anonymous):

all

OpenStudy (campbell_st):

if you are looking for the largest common factor of all the terms. identify the common factors in each term